The arithmetic mean of the 5 consecutive integers starting with 's' is 'a'. What is the arithmetic mean of 9 consecutive integers that start with s + 2?
Answer by richwmiller(17219)

s+s+1+(s+2)+s+3+s+4
(5s+10)/5=s+2=a
s+2+s+3+s+4+s+5+(s+6)+s+7+s+8+s+9+s+10
(9s+54)/9=s+6
a+4 or s+6